### v1

@@ Appendix E Omitted Proofs from Section 3
This confirms the numerator is non-positive. Since $w_{ij^{\prime\prime}}>w_{ij^{\prime}}$ we conclude that $\Delta_{\mu}\leq 0$.
@@
$=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{s_{i}^{j}}]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]=\operatorname{\mathbb{E}}[size(\mathcal{S}^{-})]+\operatorname{\mathbb{E}}[s_{i^{*}}^{j^{*}}]$
@@
For any adaptive policy, denote by $\mathcal{S}$ the random set of pairs $(i,j)$ that the policy inserts. Set $x_{ij}=\Pr[(i,j)\in\mathcal{S}]$. From the previous lemma, the expected mass that $\mathrm{ADAPT}$ attempts to insert is $\operatorname{\mathbb{E}}(\mu(\mathcal{S}))=\sum_{ij}\mu_{ij}x_{ij}\leq 2t$, therefore, it is a feasible solution to $\Phi(2t)$ with $\sum_{ij}w_{ij}x_{ij}\leq\Phi(2t)$. Since $w_{ij}x_{ij}$ is an upper bound on the expected profit of inserting the item choice $(i,j)$, we have that
@@
for any $c\geq 2$. Now, given an optimal solution $x$ to $\Phi(ct)$, we can construct a feasible solution $x^{\prime}$ to $\Phi(t)$ by setting $x^{\prime}_{ij}=\frac{x_{ij}}{c}$ for all $i,j$, and then increasing the value of $x^{\prime}_{ij_{\bot}}$ to ensure that $\sum_{j}x^{\prime}_{i,j}=1$ for all $i$. This results in a feasible solution to $\Phi(t)$ with $\sum_{ij}w_{ij}x^{\prime}_{ij}=\frac{1}{c}\sum_{ij}w_{ij}x_{ij}$. We conclude that
@@
Appendix F Proof of Lemma 2.1
@@
  • The value of item $i$ is set to the principal’s reward $v_{i}$.
  • The cost of choice $(i,j)$ is set to $p_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)]$, where $t_{i}$ is the contract that minimizes the expected payment, as given by the LP in Definition 2.1.
Now, consider any policy $\varphi$ for the contracts problem, and denote by $\varphi^{\prime}$ the corresponding policy for the knapsack problem. $\varphi$ specifies the choice of (agent, effort level) pairs for the contracts problem, and $\varphi^{\prime}$ specifies the corresponding (item, choice) pairs for the knapsack problem.
The profit for policy $\varphi$ is a random variable that depends on the (possibly random) choices of the policy and the random sizes of the attempted jobs. Its expectation is given by
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}-\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in S}{v_{i}}\right]-\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right].$ (9)
Expanding the transfer term, we get:
@@
$\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]$ $=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}{t_{i}(s_{i})}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]=\sum_{(i,j)}\operatorname{\mathbb{E}}\left[{\mathbb{1}_{(i,j)\in A}\cdot t_{i}(s_{i})}\right]$
$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot\operatorname{\mathbb{E}}_{s_{i}\sim F_{i}^{j}}\left[t_{i}(s_{i})\right]\stackrel{{\scriptstyle(\star\star)}}{{=}}\sum_{(i,j)}\operatorname{\mathbb{E}}\left[\mathbb{1}_{(i,j)\in A}\right]\cdot p_{i}^{j}$
$=\operatorname{\mathbb{E}}\left[\sum_{(i,j)}\mathbb{1}_{(i,j)\in A}\cdot p_{i}^{j}\right]=\operatorname{\mathbb{E}}\left[\sum_{(i,j)\in A}p_{i}^{j}\right],$
where equality $(\star)$ is due to the independence between $\mathbb{1}_{(i,j)\in A}$ and $s_{i}^{j}$, and equality $(\star\star)$ is due to the definition of $p_{i}^{j}$. Plugging this back in Eq. 9 yields:

### v2

@@ Appendix D Omitted Proofs from Section 3
This confirms the numerator is non-positive. Since $w_{ij^{\prime\prime}}>w_{ij^{\prime}}$ we conclude that $\Delta_{\mu}\leq 0$. ∎
@@
$=\operatorname{\mathbb{E}}[\sum_{(i,j)\in\mathcal{S}^{-}}{s_{i}^{j}}]+\operatorname{\mathbb{E}}[\min\{t,s_{i^{*}}^{j^{*}}\}]=\operatorname{\mathbb{E}}[size(\mathcal{S}^{-})]+\operatorname{\mathbb{E}}[\min\{t,s_{i^{*}}^{j^{*}}\}]$
@@
For any adaptive policy and capacity $t$, denote by $\mathcal{S}$ the random set of pairs $(i,j)$ that the policy inserts. Set $x_{ij}=\Pr[(i,j)\in\mathcal{S}]$. From the previous lemma, the expected mass that $\mathrm{ADAPT}$ attempts to insert is $\operatorname{\mathbb{E}}(\mu(\mathcal{S}))=\sum_{ij}\mu_{ij}x_{ij}\leq 2t$, therefore, it is a feasible solution to $\Phi(2t)$ with $\sum_{ij}w_{ij}x_{ij}\leq\Phi(2t)$. Since $w_{ij}x_{ij}$ is an upper bound on the expected utility of inserting the item choice $(i,j)$, we have that
@@
for any $c\geq 2$.
Now, given an optimal solution $x$ to $\Phi(ct)$, we can construct a feasible solution $x^{\prime}$ to $\Phi(t)$ by setting $x^{\prime}_{ij}=\frac{x_{ij}}{c}$ for all $i,j$, and then increasing the value of $x^{\prime}_{ij_{\bot}}$ to ensure that $\sum_{j}x^{\prime}_{i,j}=1$ for all $i$. This results in a feasible solution to $\Phi(t)$ with $\sum_{ij}w_{ij}x^{\prime}_{ij}=\frac{1}{c}\sum_{ij}w_{ij}x_{ij}$. We conclude that
@@
D.3 LP Overconfidence: An $\Omega(\alpha)$ Overestimation of the True Utility
In this section, we provide an instance of the MSKC problem where the value of the LP introduced in Eq. (1) is larger by a factor $\Omega(\alpha)$ from the obtainable value of any policy. This is in stark contrast to the case without costs, where Dean et al. [10] show a non-adaptive policy whose utility is a constant factor away from the value of the analogous LP. As our non-adaptive policy in Section 3.2 gives $O(\alpha)$-approximation to the value of the LP in Eq. (1), this approximation is the best possible with respect to the value of the LP.
In our instance, we have an infinite number of identical items with value $1$ and cost $1-\epsilon$. The size-distribution of the items is
By Definition 2.1, the value of the IOR parameter is $\alpha=\frac{1-\epsilon}{\epsilon}=O(1/\epsilon)$. For every item $i$, we have $\mu_{i}=(1-\epsilon)\cdot\epsilon^{n}+\epsilon\cdot 1=\epsilon^{n}+\epsilon-\epsilon^{n+1}\leq 2\epsilon$ and $w_{i}=\epsilon$. Setting $x_{i}=1$ for $\ell=\lfloor\frac{1}{2\epsilon}\rfloor$ items, and for an additional item $i^{\prime}$ set $x_{i^{\prime}}=\frac{1}{2\epsilon}$-$\ell$, implies that
$\sum_{i}x_{i}\cdot\mu_{i}\leq 2\epsilon\cdot\sum_{i}x_{i}=2\epsilon\cdot\frac{1}{2\epsilon}=1,$
which is a feasible assignment to the LP. Therefore, $\Phi(1)\geq\frac{1}{2\epsilon}\cdot\epsilon=1/2.$ We next bound the utility of any policy for this instance.
For any policy, after inserting the first item, we have that any item fits with probability at most $1-\epsilon$; thus, we have that the benefit from trying to insert any item after the first one is at most $(1-\epsilon)v-p=0.$ Therefore, the policy that only inserts one item is optimal, and it obtains a utility of $\epsilon\leq 2\epsilon\cdot\Phi(1)=\Phi(1)/\Omega(\alpha).$
Appendix E Appendix for Section 4
E.1 Proof of lemma 4.4
@@ Appendix E Appendix for Section 4
Let $\sigma_{A}$ be an optimal adaptive policy. Let $x_{i,j,t}$ be the probability that $\sigma_{A}$ attempts to insert $(i,j)$ at time $t$. We claim that $x$ is a feasible solution to the LP $\Psi(T)$. It is immediate that eq. 6, eq. 5 hold (by the definition of $x_{i,j,t}$, each item and choice are inserted at most once in the run of $\sigma_{A}$). It remains to show that eq. 4 holds. For any $i\in[n]$, $j\in[m]$, $t^{\prime}\in\left\{0,\ldots,T\right\}$: let $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}$ be the indicator for the event where item-choice $(i,j)$ is inserted at time $t^{\prime}$ and let $\mathbb{I}_{\left\{i,j,s\right\}}^{size}$ be the indicator for the event where item-choice $(i,j)$ realizes to size $s$; that is: $s_{i}^{j}=s$.
For any time $t$, let $(i_{t},j_{t})$ be the last item-choice pair that $\sigma_{A}$ inserts up until or at time $t$, and let $P_{t}$ be the set of all item-choices successfully inserted by $\sigma_{A}$ before inserting $(i_{t},j_{t})$. So the sum of all inserted item sizes is at most $t$.
$\sum_{i,j\in P_{t}}\sum_{t^{\prime}\leq t}\sum_{s\leq T}\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot s\leq t.$
Let $P^{\prime}_{t}=P_{t}\cup\left\{(i_{t},j_{t})\right\}$. We get $\sum_{i,j\in P^{\prime}_{t}}s_{i}^{j}=\sum_{i,j\in P_{t}}s_{i}^{j}+s_{i_{t}}^{j_{t}}$ and thus:
$\sum_{i,j\in P^{\prime}_{t}}\sum_{t^{\prime}\leq t}\sum_{s\leq T}\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot\min(s,t)\leq 2t.$
For any $t^{\prime}\leq t$ and $i\in[n],j\in[m]$: $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}=1\iff{i,j}\in P^{\prime}_{t}$ and thus we have:
$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\ \sum_{s\leq T}\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\cdot\mathbb{I}_{\left\{i,j,s\right\}}^{size}\cdot\min(s,t)\leq 2t.$
By taking expectation and using linearity of expectation we have:
@@
$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\ \sum_{s\leq T}\operatorname{\mathbb{E}}\left[\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\cdot\mathbb{I}_{\left\{i,j,s\right\}}^{size}\right]\cdot\min(s,t)=$
$\qquad=\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}\Pr\left(\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}\right)\left(\sum_{s\leq T}\Pr\left(\mathbb{I}_{\left\{i,j,s\right\}}^{size}\right)\min(s,t)\right)$
$\qquad=\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\ \mu_{i,j,t},$
where the first equality holds due to the independence between $\mathbb{I}_{\left\{i,j,t^{\prime}\right\}}^{ins}$, $\mathbb{I}_{\left\{i,j,s\right\}}^{size}$ (the events are independent as the policy inserts item-choice $(i,j)$ before seeing the realized value of $s_{i}^{j}$). Hence
$\sum_{i,j\in[n]\times[m]}\ \sum_{t^{\prime}\leq t}x_{i,j,t^{\prime}}\ \mu_{i,j,t}\leq 2t.$
We’ve shown that $x$ is indeed a feasible solution to the LP.
We now bound $E[\mathrm{ADAPT}]$. Let $W_{i,j}$ be the utility $\sigma_{A}$ gets from item-choice $i,j$ (if $\sigma_{A}$ did not attempt inserting the item-choice then the corresponding utility is 0). From linearity of expectation we get: